Shift report 20030201 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Feb 01 2003 - 07:45:04 EST

  • Next message: Apache: "Shift report 20030201 08:00-16:00"
    ------------------------------------------------------------
    Date: Sat Feb  1 07:45:04 2003, Shift: 00:00-08:00
    Supervisor: PS, with: 
    Log book#13 pages 135 - 136
    
    Shift Summary:
    ------------------------------------------------------------
    1. 3 stores during this shift -
       the last one is still lasting.
       Two first stores had ~30x10^11 ions in blue
       and ~60x10^9 ions in yellow ring. The collision (ZDC)
       rate was ~4k with low background, but in both cases
       beam was lost after about 30 minutes.
       Third store had ~30x10^11 ions in blue
       and ~30x10^9 ions in yellow with initial collision rate
       of 2k.
     
    2.Only data runs:
      7716 and 7717  MRS:6B, FS:1/15A 
      7718 and 7719  MRS:6A, FS:1/15A
      DAQ lost connection with fb3 (at 22:48 31Jan). Coudn't fix
      it, so decided to run without DCs (painful decision)
      and set FS field to 1/15.  
    
    3. NO CONNECTION WITH FB3. EXPERT NEEDED.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sat Feb 01 2003 - 07:49:57 EST